OpenStudy (anonymous):

Amy has a garden covering 900 square feet. a. If she wants to plant moss roses on 10% of the garden, how many square feet will moss roses cover? Write and solve a percent equation to solve this problem. b. If the garden occupies 25% of Amy’s yard, what is the area of her yard in square feet? Write and solve a percent equation to solve this problem. (6 points)

OpenStudy (anonymous):

a. Equation: x/900 = 1/100 x = 90 sq. ft.